Gold is the most ductile of all metals. For example, one gram of gold can be drawn into a wire 2.35km long. What is the resistance of such a wire at 20°C? (The density of gold at 20°C is 1.93 104 kg/m3.)
1
Expert's answer
2018-09-17T15:18:08-0400
V=m/ρ=(0.001 kg)/(1.93 ∙〖10〗^4 kg/m^3 ) = 5.181∙〖10〗^4 m^3
We also know that the volume of a wire (V) is given by:
V = AL
where A is the cross sectional area of the wire and L is the length
A =V/L =( 5.181∙〖10〗^4)/(2320 )= 2.233 ∙ 〖10〗^(-11) m^2
R=ρ_gold l/A=(2.44 ∙ 〖10〗^(-8) )(2320)/( 2.233 ∙ 〖10〗^(-11) )=2.53 ∙ 〖10〗^6 Ohms